The speed of sound in hydrogen at 0°C is 1280 m/s. The density of hydrogen at STP is 0.089 kg/m3. Calculate the molar heat capacities Cp and Cv of hydrogen.

Grade:upto college level

1 Answers

Aditi Chauhan
9 years ago
Sol. V = 1280 m/s, T = 0°C, foH base 2 = 0.089 kg/m^3, rR = 8.3 J/mol-k, At STP, P = 10^5 Pa, We know V base sound = √γP/fo ⇒ 1280 = √γ * 10^5/0.089 ⇒ (1280)^2 = γ * 10^5/0.089 ⇒ γ = 0.089 * (1280)^2/10^5 = 1.458 Again, C base v = R/γ – 1 = 8.3/1.458 – 1 = 18.1 J/mol-k Again, C base P/C base V = γ or C base P = γC base V = 1.458 * 18.1 = 26. 3 J/mol-k